# Discovering Nuclear Models from Symbolic Machine Learning

### v1

@@ Introduction
Recent advancements in Machine Learning (ML) have emerged as a pragmatic tool to tackle this challenge, with the expectation of fitting models that could reproduce the vast amount of available nuclear properties [10, 11, 12, 13, 7]. Nonetheless, this approach results in black boxes with obscure physical interpretability [14, 15, 16], and often its use is limited to specialized users who can have access to and run specific software. Moreover, conventional ML models rely on extensive experimental data or high-fidelity emulators, which may not always be available and are very prone to over-fitting, drastically limiting their extrapolation power [17, 18].
An alternative approach is given by symbolic ML, which aims to provide analytical expressions to describe a given observable [19]. Intriguingly, this methodology is remarkably similar to how nuclear physicists have traditionally tackled the problem of describing complex nuclear phenomena. Unlike traditional regression techniques that fit data to a pre-specified model, such as linear or polynomial functions, modern symbolic regression approaches do not assume any particular prior model. In contrast, it changes the existing functional forms in an evolutive manner [20, 21]. Because of this, discovering complex equations is now conceivable, something that could not be accomplished by evaluating every possible mathematical expression by brute force [22, 23]. These methods have been widely successfully applied in scientific research, including Cosmology [24, 25], Astronomy [26, 27], Materials science [28], Chemistry [29], Dynamics [30], and Particle physics [31, 32]. However, applying existing methods to describe nuclear physics observables poses several challenges, as most of them have not been developed to solve multi-objective problems and/or do not consider experimental uncertainties.
@@
One of the critical challenges in ML approaches is the quantification of uncertainty in predictive modeling, especially given the often limited and low-precision experimental data that exists for certain nuclear properties. To achieve an estimation of uncertainty for the MISR model, we resample different the training datasets using jackknife bootstrapping [38] to link each one of the model’s predictions to a probability distribution, from which we sample the posteriors. In addition, we include a truncation uncertainty, for a specific cutoff, given by the absolute value of the subsequent expansion term. This methodology was inspired by an analogy to how effective theories are developed [39].
@@
As a proof-of-principle of our approach, we focus on the study of nuclear binding energies and charge radii. However, our proposed method can be generally applied to describe other nuclear observables, with an arbitrary set of input variables. Binding energy, the energy required to disassemble a nucleus into its constituent protons and neutrons, provides important information to guide our understanding of atomic nuclei. This observable is arguably one of the most studied nuclear observables, with a relatively large collection of experimental data available. To train our algorithm, we used a subset of the experimental data of binding energies reported in the AME2020 mass evaluation data set [41], and experimental charge radii reported in [42]. For the charge radii —characterizing the charge distributions inside the nuclei— there is a more scarce set of experimental measurements [43]. This is particularly challenging for light isotopes, which, combined with the large variations in their radii, motivated us to focus on nuclei with $12\leq Z\leq 50$, for which reliable experimental data exists. From these data sets, we sampled uniformly $20\%$ of the nuclei for testing purposes. The cutoff on light and medium mass nuclei allows us to study the extrapolation capabilities of our algorithm and provide a test of the model’s ability to deduce complex nuclear properties from relatively small data sets. This demonstrates the broader applicability of our MIRS method, which can describe physical observables where data scarcity is common.
@@
where $\delta_{0}$ is another fitted constant [44]. More complex empirical models such as the Finite Range Droplet Model [45], and the Duflo-Zuker (DZ) [46], attempt to incorporate microscopic corrections to these macroscopic trends, refining the reproduction of nuclear binding energies across the chart of nuclides. The DZ model integrates both macroscopic and microscopic aspects of nuclear structure, featuring a wide array of terms to capture various models of nuclear interactions and nucleon configurations. The DZ model is not represented by a concise formula, but rather as a computational tool. This model is relatively complex and includes several terms to numerically account for nuclear shell effects, monopole Hamiltonian contributions, deformation energies, symmetry, and Coulomb energies.
@@
The differences between the experimental values and model predictions for various observables are shown in Fig. 2. These differences decrease as one includes more orders of the MISR expansion, illustrating how the regression convergences without a sign of overfitting to the trained region. In principle, one can keep an arbitrary number of terms in the expansion. However, the Pareto Frontier hints it is optimal to keep up to 10 terms.
Figure 2: Convergence of different observables fitting the MISR on the nuclear binding energies. The colored area illustrates the standard deviation of the residuals among the training set and the dashed line shows the metrics over the test nuclei. The LDM results are illustrated as horizontal lines for reference.
@@
In Table 1 we compare the performance of different nuclear models in terms of Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE). MAE is calculated as the average absolute difference between the predicted and experimental values, while RMSE is derived from the square root of the average of the squared differences between the predicted and experimental binding energy values. For comparison, the table includes the results from DFT calculations using different density functionals such as DD-ME2 [51], NL3* [52], HFB24 [53], and UNEDF1 [54]. Remarkably, the MISR10 model shows competitive performance, particularly when considering the fact that it is built on a small number of input variables.
@@
Inspired by the fact that our MISR model appears to be highly complementary to the DZ model, i.e. MISR exhibits better performance around closed shells, while the DZ performs better for open-shell nuclei (see Fig. 3), we decided to use to combine them as an ensemble model. Both models can be combined using a Bayesian Automatic Relevance Determination (ARD) regression [63], which can determine the weight of each model and provide a combined prediction with uncertainty. The obtained assembled model, labeled as ARD provides an MAE of 0.389 (0.411) MeV on the train (test) samples, yielding a relative improvement of around $20\%$ with respect to the best-performing HFB24 model [53] (see Table 1). This motivates us to use the ARD model to study open questions in nuclear science, such as the limits of the existence of nuclear matter. The ARD model provides a robust estimation of nucleon separation energies, which is critical for predicting the limits of stability of nuclei. Using our model uncertainties, the probability of having a bound nucleus can be estimated for any number of protons and neutrons. Proceeding with this approach, we compute the one and two nucleon separation energies and obtain the probability of having a positive central value via the Cumulative Density Probability of a Gaussian distribution [48]. The results for the limits of stability predicted by the ARD model for nuclei with $12\leq Z\leq 50$ are presented in Fig. 8. Where data are available, overall good agreement with the experiment was found.
@@
Future work will focus on extending the application of MISR to describe complementary nuclear observables using different sets of input variables. Of particular interest would be the development of the MISR method to discover the analytical forms of nuclear interactions and, more generally, interaction Hamiltonians of quantum many-body systems. This could be combined with powerful many-body methods to establish a direct link between microscopic interactions and observables. Work is ongoing to generalize the MISR approach for this purpose and for applications to other physical problems.

### v3

@@ Introduction
Recent advancements in Machine Learning (ML) have emerged as a pragmatic tool to tackle this challenge, with the expectation of fitting models that could reproduce the vast amount of available nuclear properties [10, 11, 12, 13, 14, 7, 15]. Nonetheless, this approach results in black boxes with obscure physical interpretability [16, 17, 18], and often its use is limited to specialized users who can have access to and run specific software. Moreover, conventional ML models rely on extensive experimental data or high-fidelity emulators, which may not always be available and are very prone to over-fitting, drastically limiting their extrapolation power [19, 20].
An alternative approach is given by symbolic ML, which aims to provide analytical expressions to describe a given observable [21]. Intriguingly, this methodology is remarkably similar to how nuclear physicists have traditionally tackled the problem of describing complex nuclear phenomena. Unlike traditional regression techniques that fit data to a pre-specified model, such as linear or polynomial functions, modern symbolic regression approaches do not assume any particular prior model. In contrast, it changes the existing functional forms in an evolutive manner [22, 23, 24]. Because of this, discovering complex equations is now conceivable, something that could not be accomplished by evaluating every possible mathematical expression by brute force [25, 26]. These methods have been widely successfully applied in scientific research, including Cosmology [27, 28], Astronomy [29, 30], Materials science [31], Chemistry [32], Dynamics [33], and Particle physics [34, 35]. However, applying existing methods to describe nuclear physics observables poses several challenges, as most of them have not been developed to solve multi-objective problems and/or do not consider experimental uncertainties.
@@
One of the critical challenges in ML approaches is the quantification of uncertainty in predictive modeling, especially given the often limited and low-precision experimental data that exists for certain nuclear properties. To achieve an estimation of uncertainty for the MISR model, we resample different training datasets using jackknife bootstrapping [41] to link each one of the model’s predictions to a probability distribution, from which we sample the posteriors. In addition, we include a truncation uncertainty, for a specific cutoff, given by the absolute value of the subsequent expansion term. This methodology was inspired by an analogy to how effective theories are developed [42].
@@
As a proof-of-principle of our approach, we focus on the study of nuclear binding energies and charge radii. However, our proposed method can be generally applied to describe other nuclear observables, with an arbitrary set of input variables. Binding energy, the energy required to disassemble a nucleus into its constituent protons and neutrons, provides important information to guide our understanding of atomic nuclei. This observable is arguably one of the most studied nuclear observables, with a relatively large collection of experimental data available. To train our algorithm, we used a subset of the experimental data of binding energies reported in the AME2020 mass evaluation dataset [44, 45, 46], and experimental charge radii reported in [47]. For the charge radii —characterizing the charge distributions inside the nuclei— there is a more scarce set of experimental measurements [48]. This is particularly challenging for light isotopes, which, combined with the large variations in their radii, motivated us to focus on nuclei with $12\leq Z\leq 50$, for which reliable experimental data exists. From these data sets, we sampled uniformly $20\%$ of the nuclei for testing purposes. The cutoff on light and medium mass nuclei allows us to study the extrapolation capabilities of our algorithm and provide a test of the model’s ability to deduce complex nuclear properties from relatively small data sets. This demonstrates the broader applicability of our MIRS method, which can describe physical observables where data scarcity is common.
@@
where $\delta_{0}$ is another fitted constant [49]. More complex empirical models such as the Finite Range Droplet Model (FRDM) [50], and the Duflo-Zuker (DZ) [51], attempt to incorporate microscopic corrections to these macroscopic trends, refining the reproduction of nuclear binding energies across the chart of nuclides. The DZ model integrates both macroscopic and microscopic aspects of nuclear structure, featuring a wide array of terms to capture various models of nuclear interactions and nucleon configurations. The DZ model is not represented by a concise formula, but rather as a computational tool. This model is relatively complex and includes several terms to numerically account for nuclear shell effects, monopole Hamiltonian contributions, deformation energies, symmetry, and Coulomb energies.
@@
The differences between the experimental values and model predictions for various observables are shown in Fig. 2. These differences decrease as one includes more orders of the MISR expansion, illustrating how the regression converges without a sign of overfitting to the trained region. In principle, one can keep an arbitrary number of terms in the expansion. However, the Pareto Frontier hints it is optimal to keep up to 10 terms.
Figure 2: Convergence of different observables fitting the MISR on the nuclear binding energies. The colored area illustrates the standard deviation of the residuals among the training set and the dashed line shows the mean over the test nuclei. The LDM results are illustrated as horizontal lines for reference.
@@
In Table 1 we compare the performance of different nuclear models in terms of Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE). Note that these results represent only a subset of DFT calculations and do not include the latest DFT developments. MAE is calculated as the average absolute difference between the predicted and experimental values, while RMSE is derived from the square root of the average of the squared differences between the predicted and experimental binding energy values. For comparison, the table includes the results from DFT calculations using different density functionals such as DD-ME2 [56], NL3* [57], HFB24 [58], and UNEDF1 [59]. Remarkably, the MISR10 model shows competitive performance, particularly when considering the fact that it is built on a small number of input variables.
@@
Inspired by the fact that our MISR model appears to be highly complementary to the DZ model, i.e. MISR exhibits better performance around closed shells, while the DZ performs better for open-shell nuclei (see Fig. 3), we combined them as an ensemble model. Both models can be combined using a Bayesian Automatic Relevance Determination (ARD) regression [68], which can determine the weight of each model and provide a combined prediction with uncertainty. The obtained assembled model, labeled as ARD provides an MAE of 0.389 (0.411) MeV on the train (test) samples, yielding a relative improvement of around $20\%$ with respect to the best-performing HFB24 model [58], a comparison of different models is presented in Table 3.
Table 3: Mean Absolute Error (MAE) between experiment and model for binding energy and separation energies of nuclei with $12\leq Z\leq 50$ over the complete set of nuclei (testing and training). Our results are compared with different DFT calculations and phenomenological models. See text for more details.
Model Ref. $BE$ [MeV] $S_{n}$[MeV] $S_{p}$[MeV]
MISR10 - 0.79 0.46 0.96
ARD - 0.39 0.28 0.77
Duflo-Zuker [51] 0.62 0.30 0.81
FRDM [50] 0.67 0.34 0.84
HFB24 [58] 0.55 0.48 0.91
UNEDF1 [59] 1.88 0.41 0.82
This motivates us to use the ARD model to study open questions in nuclear science, such as the limits of the existence of nuclear matter. The ARD model provides a robust estimation of nucleon separation energies, which is critical for predicting the limits of stability of nuclei. Using our model uncertainties, the probability of having a bound nucleus can be estimated for any number of protons and neutrons. Proceeding with this approach, we compute the one and two nucleon separation energies and obtain the probability of having a positive central value via the Cumulative Density Probability of a Gaussian distribution [53]. The results for the limits of stability predicted by the ARD model for nuclei with $12\leq Z\leq 50$ are presented in Fig. 8. Where data are available, overall good agreement with the experiment was found.
@@
Future work will focus on extending the application of MISR to describe complementary nuclear observables using different sets of input variables. Of particular interest would be the development of the MISR method to discover the analytical forms of nuclear interactions and, more generally, interaction Hamiltonians of quantum many-body systems. This could be combined with powerful many-body methods to establish a direct link between microscopic interactions and observables. Work is ongoing to generalize the MISR approach for this purpose, including vectors and tensor operators.
Acknowledgements
This work was supported by the Office of Nuclear Physics, U.S. Department of Energy, under grants DE-SC0021176 and DE-SC0021179. We are grateful for useful discussions and suggestions from S. Wilkins, J. Holt, A. Belley, A. Galindo, and K. Matchev for their helpful comments and insights.